OpenStudy (perl):

Find the derivative of f (x, y) = xy/ (x+ y) at the point (2, -1) in the direction of 6i + 8j i got < fx, fy> = < y^2/(x+y)^2 , x^2/(x+y)^2> , and plug in that bad boy (2,-1) then im lost , we get <1,4> x <6,8>

OpenStudy (amistre64):

\[fx = \frac{y^2}{ (x+ y)^2}\]those look good

OpenStudy (amistre64):

if the surface is not defined as a vector function but a scalar function that might throw us for a loop

OpenStudy (perl):

my calculus is 1200 pages, im scared to open it

OpenStudy (perl):

calculus book *

OpenStudy (perl):

i get the idea, they want the slope at the direction 6i + ...

OpenStudy (amistre64):

 like this?

OpenStudy (amistre64):

\[cos(t)=\frac{u.v}{|u||v|}\]

OpenStudy (amistre64):

if thats helps

OpenStudy (perl):

hmmmm , i will come back and work on it, thanks maybe paul has a solution one sec

OpenStudy (amistre64):

good luck with it :)

OpenStudy (perl):

hey i found a nifty graphing calculator www.romanlabs.com

OpenStudy (amistre64):

so you get the derivative to be <1,4> and dot that to a unit vector to get: 1u + 4u the directional vector is the unitified and subbed into this: <6,8>/10 = <3/5,4/5> 1(3/5) + 4(4/5) 3/5 + 16/5 = 19/5